An object moves across a level, frictionless surfa
Answer #1


(a)

units of axy = N

unit of x , y = m

a*m*m = kg m /s^2

a = kg m^-1 s^-2

unit of by = N

b*m = kg m s-2

b = kg s-2


=======================

(b)

from origin to (5,0)

along this line y = 0


F = 0


W = 0


towards north

x = 5


F = 7*5*y i + 4y j

F = 30y i + 4y j

work = F.r


r = yj

W = 4y^2 J = 4*3^2 J

W = 36 J


Wtotal = 36 J

================


(c)

towards north

x = 0

F = by j = 4y j

r = y j


W = F.r = 4y^2 J

W = 4*3^2 = 36 J


towards east


y = 3

F = 21x i + 12 j

r = xi


W = 21x^2 J

W = 21*5^2 = 525 J


Wtotal = 561 J